# Are We in the AI-Generated Text World Already? Quantifying and Monitoring AIGT on Social Media

### v2

@@ Abstract
Social media platforms are experiencing a growing presence of AI-Generated Texts (AIGTs). However, the misuse of AIGTs could have profound implications for public opinion, such as spreading misinformation and manipulating narratives. Despite its importance, it remains unclear how prevalent AIGTs are on social media. To address this gap, this paper aims to quantify and monitor the AIGTs on online social media platforms. We first collect a dataset (SM-D) with around $2.4M$ posts from $3$ major social media platforms: Medium, Quora, and Reddit. Then, we construct a diverse dataset (AIGTBench) to train and evaluate AIGT detectors. AIGTBench combines popular open-source datasets and our AIGT datasets generated from social media texts by $12$ LLMs, serving as a benchmark for evaluating mainstream detectors. With this setup, we identify the best-performing detector (OSM-Det). We then apply OSM-Det to SM-D to track AIGTs across social media platforms from January 2022 to October 2024, using the AI Attribution Rate (AAR) as the metric. Specifically, Medium and Quora exhibit marked increases in AAR, rising from $1.77\%$ to $37.03\%$ and $2.06\%$ to $38.95\%$, respectively. In contrast, Reddit shows slower growth, with AAR increasing from $1.31\%$ to $2.45\%$ over the same period. Our further analysis indicates that AIGTs on social media differ from human-written texts across several dimensions, including linguistic patterns, topic distributions, engagement levels, and the follower distribution of authors. We envision our analysis and findings on AIGTs in social media can shed light on future research in this domain.
@@ Introduction
The rapid development of Large Language Models (LLMs) has markedly enhanced the quality of AIGTs, enabling the use of models like GPT-3.5 [39] in daily life to produce high-quality texts, such as in academic writing [14], question-answering [22], and translation [61]. These AIGTs are often indistinguishable from Human-Written Texts (HWTs), presenting AIGT detection as a crucial yet challenging task for effective classification. On social media platforms, the use of LLMs to answer questions can contribute to the spread of misinformation [67]. Furthermore, AIGTs may be deliberately used for information manipulation or the dissemination of fake news, potentially resulting in serious societal impacts [16]. To better understand the prevalence of AIGTs on social media platforms, we aim to quantify and monitor its presence, addressing the question: On social media, are we already interacting with AI-generated texts?
Figure 1: Pipeline of quantifying AIGTs on social media.
@@
The growth in model parameters and training data has recently empowered LLMs to demonstrate exceptional language processing capabilities [66]. Since then, LLMs have gradually gained popularity, like GPT-4 [40] and Llama [56], enabling users to generate high-quality texts effortlessly. Yet, LLMs have raised concerns about potential misuse, such as fake news generation [65], academic misconduct [59], and performance degradation of training LLMs using AI content [6], making the detection of AIGTs (also known as machine-generated texts) increasingly important [11]. He et al. [17] introduce MGTBench for standardizing the evaluation of different LLMs and experimental setups within the AIGT detectors. They broadly categorize the detectors into two main types: metric-based and model-based detectors. Metric-based detectors use pre-defined metrics, such as log-likelihood, to capture the characteristics of texts [12, 36, 53]. In contrast, model-based detectors rely on trained models to distinguish between AIGTs and HWTs [51, 15, 4, 29, 20, 24]. More introduction refer to Appendix B.
@@
To train the AIGT detectors, we consider two parts of the data. First, we consider $6$ publicly available AIGT datasets and $5$ common SFT datasets to form the training dataset (see Tables A3 and A4 for dataset statistics and Appendix D for more details). Second, to increase the detector’s generalization capabilities on social media, we additionally collect data from the $3$ social media platforms ranging from January 1, 2018, to December 31, 2021. We classify this data as HWTs, given that most LLMs had not been published during this period. We also design different LLMs writing tasks to generate AIGTs that align with the characteristics of platforms (Table A1 describes the statistics details).
@@
Overall, the datasets used for training our detector and the distribution of LLM series are shown in Figure A1. This dataset includes $12$ different LLMs, with a detailed introduction provided in Appendix A. Within these datasets, the two most prevalent model series are the GPT Series, which accounts for $42.99\%$, and the Llama series, which represents $39.05\%$. GPT Series is the most widely used proprietary model and has played a pivotal role in the evolution of generative AI. As of January 2023, approximately $13M$ users interact daily with GPT-3.5 [62]. The Llama series models also have significant influences, as the report indicates that downloads of Llama models on the Hugging Face platform have nearly reached around $350M$ [35]. Therefore, these two model series are the primary focus of our dataset. During the data generation process, we notice that certain samples contain textual noise, like irrelevant or redundant information. To maintain data quality, we implement some data processing strategies (see Appendix E for details).
@@
Generalizability Experiments. To validate the generalizability of OSM-Det, we conduct experiments under various generation parameter settings, unseen LLMs, and unseen domains. Figure 2 shows that OSM-Det can maintain excellent performance on AIGTs under different generation parameters (temperature, Top-P and Top-K), with accuracy exceeding $0.990$. The generalization performance on unseen LLMs and domains stays robust. For further details, please refer to Appendix H. Thus, these results show that OSM-Det has outstanding generalizability, supporting the reliability of the results on SM-D.
@@
As shown in Table A2, OSM-Det achieves False Positive Rates (FPR) of $1.82\%$, $1.36\%$, and $1.70\%$ on Medium, Quora, and Reddit, respectively, while achieving a benchmark F1-score of $0.980$ (see Table 2). These results highlight OSM-Det’s low misclassification rate and high overall accuracy, making it a reliable choice for quantifying and monitoring AIGTs on social media.
@@
Figure 4: Comparison of Medium and Quora word frequency trends: human vs. AI preferences. (The result of Reddit is shown in Figure A5)
@@
We explore the interpretability of the OSM-Det model in the case study using two methods: Integrated Gradients [54], representing a model-dependent perspective, and Shapley Value [48], offering a model-independent perspective. Details of the two methods can be found in Section I.2.
Word-Level Analysis. In the case study of Reddit (refer to  Figures A8 and A10), words like “and”, “think” and “I” have the highest Integrated Gradients and Shapley Values, which lead model to classify texts as human-written. Meanwhile, model-specific analysis shows the words “think”, “can”, and “Online” have the lowest scores, leading to AI-generated prediction. From these observations, we note that specifying clear word-level patterns between two class is challenging because certain words, like “think”, contribute significantly to both classifications. This overlap suggests that word importance is highly context-dependent. Similar challenges are also observed on Medium and Quora (Figures A14, A16, A13 and A11).
Given this difficulty, we then turn to a different approach: a statistical analysis of high-frequency adjectives, conjunctions, and adverbs (details provided in Section I.1). These high-frequency terms are then classified into human-preferred and AI-preferred vocabularies. We then track the trends of these lexical items on SM-D.
As shown in Figures 4(a) and 4(b), the NTF of AI-preferred vocabulary on the Medium and Quora is closely aligned with the development of LLMs. Following the release of LLMs such as GPT, Llama, and the Claude series, the NTF of human-preferred vocabulary has gradually declined. Meanwhile, AI-preferred vocabulary shows an increase. These results reflect an increasing usage of LLMs for content generation by Medium and Quora platform users. In contrast, the trends on Reddit show some differences (see Figure A5). From 2022 to 2024, the NTF of human-preferred vocabulary always remains high, while the AI-preferred vocabulary consistently remains low. This indicates that Reddit users rely less on LLMs to produce content. From above, we obverse that word frequency changes closely align with the AAR trends in Figure 3.
Sentence-Level Analysis. We also conduct a sentence-level analysis using Shapley values, as Integrated Gradients are only suitable for word-level. From the case studies of Medium, Quora, and Reddit (shown in Figures A15, A12 and A9), we observe that AIGTs are characterized by their objective and standardized structures, typically beginning with a noun or pronoun and following a verb-object pattern, like “Online bullying…contributes…feelings…”. In contrast, HWTs often contain flexible sentence structures and informal expressions, as illustrated by “That being said, why not both?” and “Why can’t we restore…”. In summary, the results suggest that sentence-level patterns provide more distinctive characteristics for distinguishing AIGTs and HWTs, as LLMs may usually follow a standardized pattern to generate texts.
@@ Conclusion
    Limited coverage of LLMs: AIGTBench includes only $12$ LLMs and does not cover all LLMs released across different time periods. Although the current AIGT detectors can generalize to LLMs that are not involved in training to a certain extent [24], there may still be slight errors, which poses a potential impact on the accuracy of some results. We also note that AIGTBench exhibits a distributional bias in the number of LLM-generated texts, favoring the GPT series and Llama series models, which dominate its composition at $42.9\%$ and $39.05\%$, respectively. However, this bias is unlikely to significantly impact the analysis results, as these models are also the most widely used in real-world applications.
@@
  4. 4.
    Inability of the AIGT detector to detect texts generated by early models release before ChatGPT: During the dataset collection and training process, we only included models that were released after November 2022. This decision was made because our study specifically focuses on more powerful models such as ChatGPT. We acknowledge that this approach may lead to misclassifications for earlier models; however, this does not affect the overall experimental results or the validity of our findings.
@@ Appendix G Detailed Performance of OSM-Det
Appendix H Generalizability of OSM-Det
From Table A8, we observe that OSM-Det maintains strong detection performance across these unseen models. The lowest performance was recorded for InternVL2.5-8B, yet it still achieve an accuracy of $0.925$ and an F1-score of $0.958$. This demonstrates that OSM-Det exhibits strong generalization capability when detecting AIGTs generated by previously unseen LLMs.
Appendix I Details About the Collection of High-Frequency Words and Model Interpretation Analysis Methods
@@
Figure A1: Proportion of total sentences various LLMs, with “Others” including Alpaca 7B and Vicuna 13B.
@@
Platform # text (Human) FPR
Medium $116,303$ $1.82\%$
Quora $101,145$ $1.36\%$
Reddit $53,321$ $1.70\%$
Table A2: FPR of OSM-Det on social media platforms.
@@
Figure A5: Word frequency trends on Reddit from January 1, 2022, to July 31, 2024.
@@
Human top frequency words
‘little’, ‘small’, ‘last’, ‘able’, ‘bad’, ‘next’, ‘right’, ‘most’, ‘long’, ‘old’, ‘much’, ‘sure’, ‘great’, ‘actually’, ‘again’, ‘probably’, ‘much’, ‘very’, ‘pretty’, ‘already’, ‘since’, ‘against’, ‘yet’
AI top frequency words
‘various’, ‘significant’, ‘positive’, ‘complex’, ‘original’, ‘free’, ‘specific’, ‘unique’, ‘crucial’, ‘clear’, ‘human’, ‘personal’, ‘essential’, ‘particularly’, ‘especially’, ‘truly’, ‘instead’, ‘here’, ‘rather’, ‘additionally’, ‘despite’, ‘due to’, ‘following’
@@
Table A7: Test OSM-Det in the wild. (all datasets from HuggingFace)
Model Accuracy F1-score
Deepseek-V3 0.986 0.993
GLM-4-Flash 0.997 0.998
Gemini-1.5-Flash 0.938 0.952
Gemini-2.0-Flash 0.984 0.992
Yi-1.5-34B 0.999 0.999
InternVL2.5-8B 0.925 0.958
Dolphin3.0-Llama3.1-8B 0.996 0.998
Llama3-OpenBioLLM-8B 0.960 0.980
Xwin-LM-13B-V0.2 0.996 0.998
Table A8: Performance of OSM-Det on AIGTs generated from unseen LLMs based on social media data from AIGTBench.

### v3

@@ Abstract
Social media platforms are experiencing a growing presence of AI-Generated Texts (AIGTs). However, the misuse of AIGTs could have profound implications for public opinion, such as spreading misinformation and manipulating narratives. Despite its importance, it remains unclear how prevalent AIGTs are on social media. To address this gap, this paper aims to quantify and monitor the AIGTs on online social media platforms. We first collect a dataset (SM-D) with around $2.4M$ posts from $3$ major social media platforms: Medium, Quora, and Reddit. Then, we construct a diverse dataset (AIGTBench) to train and evaluate AIGT detectors. AIGTBench combines popular open-source datasets and our AIGT datasets generated from social media texts by $12$ LLMs, serving as a benchmark for evaluating mainstream detectors. With this setup, we identify the best-performing detector (OSM-Det). We then apply OSM-Det to SM-D to track AIGTs across social media platforms from January 2022 to October 2024, using the AI Attribution Rate (AAR) as the metric. Specifically, Medium and Quora exhibit marked increases in AAR, rising from $1.77\%$ to $37.03\%$ and $2.06\%$ to $38.95\%$, respectively. In contrast, Reddit shows slower growth, with AAR increasing from $1.31\%$ to $2.45\%$ over the same period. Our further analysis indicates that AIGTs on social media differ from human-written texts across several dimensions, including linguistic patterns, topic distributions, engagement levels, and the follower distribution of authors. We envision our analysis and findings on AIGTs in social media can shed light on future research in this domain. Our code and dataset are publicly available. [^1] Code & Dataset: https://github.com/TrustAIRLab/AIGT_on_Social_Media.
@@ Introduction
The rapid development of Large Language Models (LLMs) has markedly enhanced the quality of AIGTs, enabling the use of models like GPT-3.5 [42] in daily life to produce high-quality texts, such as in academic writing [14], question-answering [23], and translation [65]. These AIGTs are often indistinguishable from Human-Written Texts (HWTs), presenting AIGT detection as a crucial yet challenging task for effective classification. On social media platforms, the use of LLMs to answer questions can contribute to the spread of misinformation [72]. Furthermore, AIGTs may be deliberately used for information manipulation or the dissemination of fake news, potentially resulting in serious societal impacts [16]. To better understand the prevalence of AIGTs on social media platforms, we aim to quantify and monitor their presence, addressing the question: On social media, are we already interacting with AI-generated texts?
Figure 1: Pipeline for quantifying AIGTs on social media: SM-D (2.4M posts), AIGTBench (training benchmark), OSM-Det (optimal detector).
@@
The growth in model parameters and training data has recently empowered LLMs to demonstrate exceptional language processing capabilities [70]. Since then, LLMs have gradually gained popularity, like GPT-4 [43] and Llama [60], enabling users to generate high-quality texts effortlessly. Yet, LLMs exhibit multiple inherent vulnerabilities [18, 57, 29, 71] and have raised concerns about potential misuse, such as fake news generation [69], academic misconduct [63], hate speech generation [52], and performance degradation of training LLMs using AI content [6], making the detection of AIGTs (also known as machine-generated texts) increasingly important [11]. He et al. [17] introduce MGTBench for standardizing the evaluation of different LLMs and experimental setups within the AIGT detectors. They broadly categorize the detectors into two main types: metric-based and model-based detectors. Metric-based detectors use pre-defined metrics, such as log-likelihood, to capture the characteristics of texts [12, 39, 56]. In contrast, model-based detectors rely on trained models to distinguish between AIGTs and HWTs [54, 15, 4, 31, 21, 25]. More introduction refer to Appendix B.
@@
Figure 2: Proportion of total sentences various LLMs, with “Others” including Alpaca 7B and Vicuna 13B.
To train the AIGT detectors, we consider two parts of the data. First, we consider $6$ publicly available AIGT datasets and $5$ common SFT datasets to form the training dataset (see Tables A3 and A4 for dataset statistics and Appendix D for more details). Second, to increase the detector’s generalization capabilities on social media, we additionally collect data from the $3$ social media platforms ranging from January 1, 2018, to December 31, 2021, as verified by an ablation study demonstrating that these new subsets fill a gap that older benchmarks missed (see Appendix J). We classify this data as HWTs, given that most LLMs had not been published during this period. We also design different LLMs writing tasks to generate AIGTs that align with the characteristics of platforms (Table A1 describes the statistics details).
@@
Overall, the datasets used for training our detector and the distribution of LLM series are shown in Figure 2. This dataset includes $12$ different LLMs, with a detailed introduction provided in Appendix A. Within these datasets, the two most prevalent model series are the GPT Series, which accounts for $42.99\%$, and the Llama series, which represents $39.05\%$. GPT Series is the most widely used proprietary model and has played a pivotal role in the evolution of generative AI. As of January 2023, approximately $13M$ users interact daily with GPT-3.5 [66]. The Llama series models also have significant influences, as the report indicates that downloads of Llama models on the Hugging Face platform have nearly reached around $350M$ [38]. Therefore, these two model series are the primary focus of our dataset. During the data generation process, we notice that certain samples contain textual noise, like irrelevant or redundant information. To maintain data quality, we implement some data processing strategies (see Appendix E for details).
@@
Platform # text (Human) FPR
Medium $116,303$ $1.82\%$
Quora $101,145$ $1.36\%$
Reddit $53,321$ $1.70\%$
Table 3: FPR of OSM-Det on social media platforms.
Generalizability of OSM-Det
@@
Model Accuracy F1-score
Deepseek-V3 0.986 0.993
GLM-4-Flash 0.997 0.998
Gemini-1.5-Flash 0.938 0.952
Gemini-2.0-Flash 0.984 0.992
Yi-1.5-34B 0.999 0.999
InternVL2.5-8B 0.925 0.958
Dolphin3.0-Llama3.1-8B 0.996 0.998
Llama3-OpenBioLLM-8B 0.960 0.980
Xwin-LM-13B-V0.2 0.996 0.998
Table 4: Performance of OSM-Det on AIGTs generated from unseen LLMs based on social media data from AIGTBench.
From Table 4, we observe that OSM-Det maintains strong detection performance across these unseen models. The lowest performance was recorded for InternVL2.5-8B, yet it still achieve an accuracy of $0.925$ and an F1-score of $0.958$. This demonstrates that OSM-Det exhibits strong generalization capability when detecting AIGTs generated by previously unseen LLMs.
@@
As shown in Table 3, OSM-Det achieves False Positive Rates (FPR) of $1.82\%$, $1.36\%$, and $1.70\%$ on Medium, Quora, and Reddit, respectively, while achieving a benchmark F1-score of $0.980$ (see Table 2). These results highlight OSM-Det’s low misclassification rate and high overall accuracy, making it a reliable choice for quantifying and monitoring AIGTs on social media.
@@
(c) Word frequency trends on Reddit from January 1, 2022, to July 31, 2024.
Figure 5: Comparison of Medium and Quora word frequency trends: human vs. AI preferences.
@@
We explore the interpretability of the OSM-Det model in the case study using two methods: Integrated Gradients [58], representing a model-dependent perspective, and Shapley Value [32], offering a model-independent perspective. Details of the two methods can be found in Section H.2.
Word-Level Analysis. In the case study of Reddit (refer to  Figures A6 and A8), words like “and”, “think” and “I” have the highest Integrated Gradients and Shapley Values, which lead model to classify texts as human-written. Meanwhile, model-specific analysis shows the words “think”, “can”, and “Online” have the lowest scores, leading to AI-generated prediction. From these observations, we note that specifying clear word-level patterns between two class is challenging because certain words, like “think”, contribute significantly to both classifications. This overlap suggests that word importance is highly context-dependent. Similar challenges are also observed on Medium and Quora (Figures A12, A14, A11 and A9).
Given this difficulty, we then turn to a different approach: a statistical analysis of high-frequency adjectives, conjunctions, and adverbs (details provided in Section H.1). These high-frequency terms are then classified into human-preferred and AI-preferred vocabularies. We then track the trends of these lexical items on SM-D.
As shown in Figures 5(a) and 5(b), the NTF of AI-preferred vocabulary on the Medium and Quora is closely aligned with the development of LLMs. Following the release of LLMs such as GPT, Llama, and the Claude series, the NTF of human-preferred vocabulary has gradually declined. Meanwhile, AI-preferred vocabulary shows an increase. These results reflect an increasing usage of LLMs for content generation by Medium and Quora platform users. In contrast, the trends on Reddit show some differences. From 2022 to 2024, the NTF of human-preferred vocabulary always remains high, while the AI-preferred vocabulary consistently remains low. This indicates that Reddit users rely less on LLMs to produce content. From above, we obverse that word frequency changes closely align with the AAR trends in Figure 4.
Sentence-Level Analysis. We also conduct a sentence-level analysis using Shapley values, as Integrated Gradients are only suitable for word-level. From the case studies of Medium, Quora, and Reddit (shown in Figures A13, A10 and A7), we observe that AIGTs are characterized by their objective and standardized structures, typically beginning with a noun or pronoun and following a verb-object pattern, like “Online bullying…contributes…feelings…”. In contrast, HWTs often contain flexible sentence structures and informal expressions, as illustrated by “That being said, why not both?” and “Why can’t we restore…”. In summary, the results suggest that sentence-level patterns provide more distinctive characteristics for distinguishing AIGTs and HWTs, as LLMs may usually follow a standardized pattern to generate texts.
@@ Conclusion
    Limited coverage of LLMs: AIGTBench includes only $12$ LLMs and does not cover all LLMs released across different time periods. We only included models released after November 2022. This decision was made because our study specifically focuses on more powerful models, such as ChatGPT, which may lead to misclassifications for earlier models. In addition, the data set shows distributional bias favoring the GPT series $42.9\%$ and the Llama series $39.05\%$ models. While current AIGT detectors can generalize to unseen LLMs to some extent [25], these coverage limitations may introduce slight errors and pose potential impacts on the accuracy of some results. However, these biases are unlikely to significantly impact the analysis results, as these models are also the most widely used in real-world applications.
@@
Acknowledgement
We would like to thank all anonymous reviewers, Area Chair, and Program Chair for their insightful comments and constructive suggestions. We are also grateful to Dr. Yun Shen for his valuable guidance and feedback in this work.
@@ Appendix G Detailed Performance of OSM-Det
Appendix H Details About the Collection of High-Frequency Words and Model Interpretation Analysis Methods
@@
Appendix I Deeper Analysis of the Result
Regarding the results of “Evaluation on Social Media Platforms”, the AAR observed on Medium, Quora, and Reddit show divergent trends in AI adoption. Medium’s consistently high AAR suggests that its longer-form, polished content format may be particularly conducive to AI-assisted creation. The early and substantial adoption may indicate that Medium’s creators see AI tools as valuable for improving article efficiency. While Medium’s policy permits AI-assisted content with disclosure, our analysis (based on syntactic patterns and keyword detection) finds that just 0.81% of authors compiled. This suggests that many users may not be following the platform’s guidelines. On the other hand, Quora’s adoption pattern tends to fluctuate. The platform experienced a temporary surge after introducing the Poe (AI tools), but this trend did not persist, possibly indicating a misalignment between AI capabilities and community expectations. In contrast, Reddit maintains a low AAR, strongly resisting AIGT despite the platform’s text-centric nature. This phenomenon may stem from Reddit’s vertically organized subcommunities (subreddits), where members reinforce identity through shared jargon and memes, content that current LLMs struggle to generate. On top of that, active community moderation behaviors serve as a self-purifying mechanism, such as downvoting "overly fluent" or suspicious content. Thus, the platform’s strong community governance and inherent culture pose barriers to AI content adoption.
Regarding the underlying reasons for “significantly higher AAR in technical fields (including Technology and Software Development)”, One possible explanation is that technical content tends to be highly structured, which fits well with the generation patterns of LLMs. Another factor is that technical terminology inherently possesses objectivity, with readers typically focusing more on information accuracy rather than storytelling, making AIGTs easier to accept. Moreover, technical communities tend to view AI tools as “advanced productivity enhancers” rather than ethical threats, further normalizing the use of AIGTs in technical subjects.
The phenomenon of “AIGTs receiving lower engagement than HWTs but with limited disparity” may be related to cognitive psychology factors. We speculate that some users possibly rely on heuristic judgments (such as language fluency and information density) to evaluate content, making it difficult for them to clearly distinguish between high-quality AIGTs and HWTs. This explains why some AIGT still achieves fundamental interaction measures. However, when content involves subjective perspectives, users tend to activate deeper analytical processes, and AIGTs often lack personal experiential grounding. This flattening of affective expression triggers user vigilance, consequently diminishing interaction behaviors.
When it comes to why “low-follower authors tend to rely more on AI”, one likely explanation is that some authors with smaller followers (0-1K) are often less experienced in writing quality content, leading them to prioritize using AI to sustain their content output. Meanwhile, high-follower authors (>5K) often consciously limit their use of AI. By doing so, they can preserve their authentic human writing style, which helps reinforce their expert credibility and protects their relationship with their audience.
We have the following recommendations for platform operators, content creators, and readers:
  • For platform operators: AI-generated content presents dual challenges of content authenticity and user trust. Our research shows that although platforms have established AI content disclosure policies, compliance rates fall far below expectations. Platforms may choose to implement defensive strategies, such as AI detection and content screening, or integrative strategies incorporating AI as a creative aid. Looking ahead, we foresee that AI content labeling will become standardized, alliances will form among platforms to harmonize policies, and the relationship between AI and human creation will evolve from substitution to symbiosis. This requires platforms to develop mechanisms that foster such an ecosystem proactively.
  • For content creators: We recommend establishing clear guidelines for AI usage: AI tools should serve as assistants, primarily supporting essential tasks such as grammar checks and formatting adjustments. The core aspects of creation, central arguments, narrative logic, and emotional expression, must remain human-driven. Besides, if human-original content is mistakenly flagged as AIGT, we encourage authors to appeal and request formal disclosure of the decision basis. This demand for transparency safeguards creators’ rights and drives platforms to refine their AIGT detection algorithms, leading to fairer content evaluation mechanisms.
  • For readers: We recommend cultivating critical thinking when engaging with AI content. For instance, check for AI assistance disclosures at the beginning of articles, analyze texts to learn general AI-generated features, and maintain skepticism toward factual claims. AI may provide incorrect answers to complex questions. Furthermore, when encountering uncertain content, readers can use accessible tools (e.g., GPTZero) to aid judgment. We also urge readers to report unlabeled AI content to platform administrators, helping maintain the community’s integrity.
@@ Appendix I Deeper Analysis of the Result
Appendix J Ablation Study on Social Media Data
To validate the necessity of the social media data collection for AIGTBench, we conducted an ablation study comparing OSM-Det performance with and without the social media training subsets on AIGTBench. As shown in Table A2, OSM-Det trained only on open-source datasets exhibits poor performance when evaluated on social media test sets, particularly on Quora and Reddit platforms. This performance gap demonstrates that traditional benchmarks fail to capture the linguistic patterns and stylistic variations inherent in social media content. The new collected datasets created address this issue, improving model robustness.
Table A2: Performance comparison of OSM-Det trained on AIGTBench without social media data.
Platform Accuracy F1-score
Medium 0.975 0.967
Quora 0.684 0.678
Reddit 0.631 0.608
@@
Human top frequency words ‘little’, ‘small’, ‘last’, ‘able’, ‘bad’, ‘next’, ‘right’, ‘most’, ‘long’, ‘old’, ‘much’, ‘sure’, ‘great’, ‘actually’, ‘again’, ‘probably’, ‘much’, ‘very’, ‘pretty’, ‘already’, ‘since’, ‘against’, ‘yet’
AI top frequency words ‘various’, ‘significant’, ‘positive’, ‘complex’, ‘original’, ‘free’, ‘specific’, ‘unique’, ‘crucial’, ‘clear’, ‘human’, ‘personal’, ‘essential’, ‘particularly’, ‘especially’, ‘truly’, ‘instead’, ‘here’, ‘rather’, ‘additionally’, ‘despite’, ‘due to’, ‘following’
@@
Table A7: Test OSM-Det in the wild (all datasets from HuggingFace).